OpenStudy (anonymous):

Solve x2 - 7x = -13. x equals quantity of negative 7 plus or minus I square root of 6 all over 2 x equals quantity of negative 7 plus or minus I square root of 3 all over 2 x equals quantity of 7 plus or minus I square root of 6 all over 2 x equals quantity of 7 plus or minus I square root of 3 all over 2

OpenStudy (solomonzelman):

\(\large\color{blue}{ \rm x^2 - 7x = -13}\) \(\large\color{blue}{ \rm x^2 - 7x +13=0}\) \(\LARGE\color{blue}{ \rm \frac{-(-7)±\sqrt{(-7)^2-4(1)(13)} }{2(1)} }\) \(\LARGE\color{blue}{ \rm \frac{7±\sqrt{49-52} }{2}}\) \(\LARGE\color{blue}{ \rm \frac{7±\sqrt{-3} }{2}}\)

OpenStudy (solomonzelman):

You will only get imaginary solutions. No real solutions.

OpenStudy (anonymous):

thank you I appreciate your help

OpenStudy (solomonzelman):

my pleasure !